OpenStudy (anonymous):

what is 2/5 divided by 3 1/2

OpenStudy (anonymous):

\[\frac{2}{5} \div 3\frac{1}{2}\]\[\frac{2}{5} \div \frac{7}{2}\]\[\frac{2}{5} \times \frac{2}{7} = \frac{2 \times 2}{5 \times 7} = \frac{4}{35}\]